# Leptogenesis in Realistic Flipped SU(5)

### v1

@@ 2 Flipped SU(5) and the Fermion masses
Superfields SM decomposition; Role in the model
@@
Chiral spectrum $\bar{f}=(\bar{\bf 5},+\frac{3}{2})$ ($u^{c},L$)
@@
$h=({\bf 5},+1)$ $(D,h_{d})$; Generate Dirac fermion masses
$\bar{h}=(\overline{\bf 5},-1)$ $(\bar{D},h_{u})$; for leptons and up,down quarks
Higgs multiplets $H=(\overline{\bf 10},{+\frac{1}{2}})$ ( $Q_{H},d_{H}^{c},\nu_{H}^{c}$); Generate $\nu^{c}$ mass and
$\bar{H}=({\bf 10},-\frac{1}{2})$ ( $\bar{Q}_{H},\bar{d}_{H}^{c},\bar{\nu}_{H}^{c}$); trigger $U(1)_{B-L}$ breaking
@@
${\cal W}_{d}$ $=$ $(Y_{d})^{*}_{ij}\,({\bf 10},{-\frac{1}{2}})_{i}\cdot({\bf 10},{-\frac{1}{2}})_{j}\cdot({\bf 5},{1})\;\to\;(Y_{d})^{*}_{ij}\,Q_{i}\,d^{c}_{j}\,h_{d}\,,$
${\cal W}_{u}$ $=$ $(Y_{u})^{*}_{ij}\,({\bf 10},{-\frac{1}{2}})_{i}\cdot(\bar{{\bf 5}},{\frac{3}{2}})_{j}\cdot(\bar{{\bf 5}},{-1})\;\to\;(Y_{u})^{*}_{ij}\,[Q_{i}\,u^{c}_{j}+\nu^{c}_{i}\,L_{j}]\,h_{u}\,,$
${\cal W}_{l}$ $=$ $(Y_{l})^{*}_{ij}\,({\bf 1},{-\frac{5}{2}})_{j}\cdot(\bar{{\bf 5}},{\frac{3}{2}})_{i}\cdot({\bf 5},1)\;\rightarrow\;(Y_{l})^{*}_{ij}\,e^{c}_{j}\,L_{i}\,h_{d}\,.$ (1)
@@ 3 The flavour pattern
We first check in the simplified case with vanishing $\delta_{M}$ and $\kappa$. Using the seesaw formula, we obtain $M_{\nu}^{0}$
$M_{\nu}^{0}=\frac{v_{u}^{2}}{M}\left(\begin{array}[]{ccc}y_{u}^{2}&0&0\\ 0&y_{c}^{2}&0\\ 0&0&y_{t}^{2}\\ \end{array}\right)-\frac{v_{u}^{2}}{M}\left(\begin{array}[]{ccc}y_{u}^{2}U_{R,31}^{2}&y_{u}y_{c}U_{R,31}U_{R,32}&y_{u}y_{t}U_{R,31}U_{R,33}\\ y_{u}y_{c}U_{R,31}U_{R,32}&y_{c}^{2}U_{R,32}^{2}&y_{c}y_{t}U_{R,32}U_{R,33}\\ y_{u}y_{t}U_{R,31}U_{R,33}&y_{c}y_{t}U_{R,32}U_{R,33}&y_{t}^{2}U_{R,33}^{2}\\ \end{array}\right)$ (27)
where $U_{R,ij}$ is the $(i,j)$ entry of $U_{R}$. This matrix predicts a vanishing neutrino mass and two heavier ones proportional to $y_{c}^{2}$ and $y_{t}^{2}$, respectively. The heavier masses are too hierarchical and conflict with neutrino oscillation data. To reproduce the correct mass hierarchy for light neutrinos, we must assume $U_{R,31},U_{R,32}\simeq{\cal O}(y_{c}/y_{t})$, i.e.,
@@ 4 Leptogenesis
$\tilde{Y}=U_{l}^{\dagger}Y_{\nu}U_{R}=V_{\rm PMNS}U_{\nu}^{\dagger}\hat{Y}_{u}U_{R}\,,$ (52)
@@
$g(x)$ $=$ $\sqrt{x}\left[\frac{1}{1-x}+1-(x+1)\log\left(\frac{x+1}{x}\right)\right]\,,$
@@
Since we are focusing on the nearly-degenerate case, then $M_{2}-M_{1}\ll M_{1}+M_{2}$. It is worth mentioning that the above formula does not hold in the resonant case $M_{i}-M_{j}\lesssim\Gamma_{i}$, i.e., $\delta_{M}\lesssim(\tilde{Y}^{\dagger}\tilde{Y})_{ii}/(4\pi)$ [18, 19]. Fortunately, as we are going to discuss below, in the parameter space we considered we have always $M_{2}-M_{1}\gg\Gamma_{1,2}$, and thus we are safe to use just the density matrix formalism with the CP source in Eq. (51). In Fig. 2 we show plots of the CP asymmetry $\epsilon^{(i)}_{\alpha\alpha}$ for the decay $N_{i}\to h_{u}L_{\alpha}$ and its CP conjugate process.
@@ Acknowledgements
This work is supported by the STFC Consolidated Grant ST/L000296/1 and the European Union’s Horizon 2020 Research and Innovation programme under Marie Sklodowska-Curie grant agreement HIDDeN European ITN project (H2020-MSCA-ITN-2019//860881-HIDDeN) (S.F.K.), and National Natural Science Foundation of China (NSFC) under Grants Nos. 12205064, 12347103 and Zhejiang Provincial Natural Science Foundation of China under Grant No. LDQ24A050002 (Y.L.Z.). L.M. is supported by a Ramoǹ y Cajal contract with Ref. RYC2020-029248-I, the Spanish National Grant PID2022-137268NA-C55 and Generalitat Valenciana through the grant CIPROM/22/69. S.F.K., G.L., and Y.L.Z. are grateful to the Mainz Institute for Theoretical Physics (MITP) of the Cluster of Excellence PRISMA+ (Project ID 390831469), for its hospitality and support.

### v2

@@ 2 Flipped SU(5) and the Fermion masses
Superfields SM decomposition Role in the model
@@
Matters $\bar{f}=(\bar{\bf 5},+\frac{3}{2})$ ($u^{c},L$) SM matters & RH neutrinos
@@
$h=({\bf 5},+1)$ $(D,h_{d})$ Generate Dirac fermion masses
$\bar{h}=(\overline{\bf 5},-1)$ $(\bar{D},h_{u})$ for leptons and up,down quarks
Higgses $H=({\bf 10},-\frac{1}{2})$ ( $Q_{H},d_{H}^{c},\nu_{H}^{c}$) Generate $\nu^{c}$ mass and
$\bar{H}=(\overline{\bf 10},{+\frac{1}{2}})$ ( $\bar{Q}_{H},\bar{d}_{H}^{c},\bar{\nu}_{H}^{c}$) trigger $U(1)_{B-L}$ breaking
@@
${\cal W}_{d}$ $=$ $(Y_{d})^{*}_{ij}\,F_{i}\,F_{j}\,h\;\to\;(Y_{d})^{*}_{ij}\,Q_{i}\,d^{c}_{j}\,h_{d}\,,$
${\cal W}_{u}$ $=$ $(Y_{u})^{*}_{ij}\,F_{i}\,\bar{f}_{j}\,\bar{h}\,\;\to\;(Y_{u})^{*}_{ij}\,[Q_{i}\,u^{c}_{j}+\nu^{c}_{i}\,L_{j}]\,h_{u}\,,$
${\cal W}_{l}$ $=$ $(Y_{l})^{*}_{ij}\;e^{c}_{j}\,\bar{f}_{i}\,h\,\;\;\to\;(Y_{l})^{*}_{ij}\,e^{c}_{j}\,L_{i}\,h_{d}\,.$ (1)
@@
We further check if there would be additional dim-4 superpotential terms (i.e., dim-5 operators) contributing to fermion masses. Those invariant under $SU(5)$ and $U(1)_{\chi}$ are
$(F_{i}\bar{f}_{j})_{\bf 5}(HH)_{\bar{\bf 5}_{S}},\quad(F_{i}\bar{f}_{j})_{\bf 5,45}(\bar{h}\Sigma)_{\bf 5,45},\quad(F_{i}F_{j})_{\bf 5,45}(h\Sigma)_{\bf 5,45}.$ (5)
The first operator vanishes at the VEV of $H$ and thus has no contribution to fermion masses. The second and third ones, after $\Sigma$ gains the VEV at the GUT scale and $h$ gains VEV at the EW scale, give contributions to up-type quark, neutrino masses and down-type quark, respectively. These terms can be forbidden by introducing a parity symmetry [4], here more precisely, $\Sigma\to-\Sigma$.
@@ 3 The flavour pattern
We first check in the simplified case with vanishing $\delta_{M}$ and $\kappa$. Using the seesaw formula $M_{\nu}^{0}=Y_{\nu}(M_{R}^{-1})^{0}Y_{\nu}^{T}v_{u}^{2}$ where $(M_{R}^{-1})^{0}=M_{R}^{-1}\big{|}_{\delta_{M}=\kappa=0}$ is denoted and $Y_{\nu}=\hat{Y}_{u}$ is considered, we obtain
$M_{\nu}^{0}\!=\!\!\frac{v_{u}^{2}}{M}\left(\begin{array}[]{ccc}y_{u}^{2}(U_{11}^{2}+U_{12}^{2})&\!\!y_{u}y_{c}(U_{11}U_{21}+U_{12}U_{22})\!\!&\!\!y_{u}y_{t}(U_{11}U_{31}+U_{12}U_{32})\!\!\\ \!\!y_{u}y_{c}(U_{11}U_{21}+U_{12}U_{22})\!\!&y_{c}^{2}(U_{21}^{2}+U_{22}^{2})&\!\!y_{c}y_{t}(U_{21}U_{31}+U_{22}U_{32})\!\!\\ \!\!y_{u}y_{t}(U_{11}U_{31}+U_{12}U_{32})\!\!&\!\!y_{u}y_{t}(U_{21}U_{31}+U_{22}U_{32})\!\!&y_{t}^{2}(U_{31}^{2}+U_{32}^{2})\\ \end{array}\right)\,,$ (25)
where $U_{ij}$ is the abbreviation of the $(i,j)$ entry of $U_{R}$, i.e., $U_{R,ij}$. Since $\det M_{\nu}^{0}\propto\det(M_{R}^{-1})^{0}=0$, there will always be one eigenvalue of $M_{\nu}$ vanishing. For the two non-vanishing masses, naively, if all entries of $U_{R}$ are assumed to be domestically distributed, one can predict them proportional to $y_{c}^{2}$ and $y_{t}^{2}$, respectively. They are too hierarchical and conflict with neutrino oscillation data. To solve this problem, we will assume a special structure in $M_{\nu}^{0}$ as below. With the help of orthogonal condition $U_{R,i1}U_{R,j1}+U_{R,i2}U_{R,j2}=\delta_{ij}-U_{R,i3}U_{R,j3}$, we re-write $M_{\nu}^{0}$ in the form
$M_{\nu}^{0}=\frac{v_{u}^{2}}{M}\left(\begin{array}[]{ccc}y_{u}^{2}&0&0\\ 0&y_{c}^{2}&0\\ 0&0&y_{t}^{2}\\ \end{array}\right)-\frac{v_{u}^{2}}{M}\left(\begin{array}[]{ccc}y_{u}^{2}U_{R,13}^{2}&y_{u}y_{c}U_{R,13}U_{R,23}&y_{u}y_{t}U_{R,13}U_{R,33}\\ y_{u}y_{c}U_{R,13}U_{R,23}&y_{c}^{2}U_{R,23}^{2}&y_{c}y_{t}U_{R,23}U_{R,33}\\ y_{u}y_{t}U_{R,13}U_{R,33}&y_{c}y_{t}U_{R,23}U_{R,33}&y_{t}^{2}U_{R,33}^{2}\\ \end{array}\right)\,.$ (32)
To reproduce the correct mass hierarchy for light neutrinos, we must assume $U_{R,31},U_{R,32}\simeq{\cal O}(y_{c}/y_{t})$, i.e.,
@@ 4 Leptogenesis
$\tilde{Y}=U_{l}^{\dagger}Y_{\nu}U_{R}=U_{\rm PMNS}U_{\nu}^{\dagger}\hat{Y}_{u}U_{R}\,,$ (57)
@@
$g(x)$ $=$ $\sqrt{x}\left[(x+1)\log\left(\frac{x+1}{x}\right)-\frac{2-x}{1-x}\right]\,,$
@@
Since we are focusing on the nearly-degenerate case, then $M_{2}-M_{1}\ll M_{1}+M_{2}$. It is worth mentioning that the above formula does not hold in the resonant case $M_{i}-M_{j}\lesssim\Gamma_{i}$, i.e., $\delta_{M}\lesssim(\tilde{Y}^{\dagger}\tilde{Y})_{ii}/(4\pi)$ [18, 19]. Fortunately, as we are going to discuss below, in the parameter space we considered we have always $M_{2}-M_{1}\gg\Gamma_{1,2}$, and thus we are safe to use just the density matrix formalism with the CP source in Eq. (56). In Fig. 2 we show plots of the CP asymmetry $\epsilon^{(i)}_{\alpha\alpha}$ for the decay $N_{i}\to h_{u}L_{\alpha}$ and its CP conjugate process. Considering we have assumed a real matrix $U_{R}$, it is straightforward to show that $\tilde{Y}^{\dagger}\tilde{Y}=U_{R}^{T}\hat{Y}_{u}^{2}U_{R}$ is real and symmetric. We can further derive the analytical formula of $\epsilon_{\alpha\alpha}^{(i)}$ approximately in the nearly-degenerate but non-resonant regime for RHN neutrino masses, i.e., $\Gamma\ll M_{2}-M_{1}\ll M_{2}+M_{1}$. And they are given by
$\epsilon_{\alpha\alpha}^{(1)}\simeq\frac{(\tilde{Y}^{\dagger}\tilde{Y})_{12}}{8\pi(\tilde{Y}^{\dagger}\tilde{Y})_{11}}\,{\rm Im}\left(\tilde{Y}_{\alpha 2}\tilde{Y}^{*}_{\alpha 1}\right)\,\frac{M_{2}+M_{1}}{M_{2}-M_{1}}\,,\quad\epsilon_{\alpha\alpha}^{(2)}\simeq\epsilon_{\alpha\alpha}^{(1)}\frac{(\tilde{Y}^{\dagger}\tilde{Y})_{11}}{(\tilde{Y}^{\dagger}\tilde{Y})_{22}}\,.$ (59)
Recall Eq. (57) that the only CP-violating phases in $\tilde{Y}$ is from the PMNS matrix. Thus, we obtain a direct connection between CP violation in the heavy RHN neutrino decay and that in light neutrino experiments. However, this result is based on the assumption of a real $U_{R}$, which is crucial in deriving the flavour pattern as well as the correlation between light and heavy neutrino masses in the last section. Without this assumption, i.e., adding some phases to $U_{R}$, we will lose the prediction.
@@ Acknowledgements
This work is supported by the STFC Consolidated Grant ST/L000296/1 and the European Union’s Horizon 2020 Research and Innovation programme under Marie Sklodowska-Curie grant agreement HIDDeN European ITN project (H2020-MSCA-ITN-2019//860881-HIDDeN) (S.F.K.), and National Natural Science Foundation of China (NSFC) under Grants Nos. 12205064, 12347103 and Zhejiang Provincial Natural Science Foundation of China under Grant No. LDQ24A050002 (Y.L.Z.). L.M. is supported by the Spanish National Grant PID2022-137268NA-C55 and Generalitat Valenciana through the grant CIPROM/22/69. S.F.K., G.L., and Y.L.Z. are grateful to the Mainz Institute for Theoretical Physics (MITP) of the Cluster of Excellence PRISMA+ (Project ID 390831469), for its hospitality and support.